# Open-FinLLMs: Open Multimodal Large Language Models for Financial Applications

### v2

@@ 3.2 Performance of FinLLaMA-Instruct
Table 10 summarizes the datasets and tasks used to evaluate instruction fine-tuned models, aligned with FinTral’s (Bhatia et al., 2024) settings for consistency. The evaluation spans 6 tasks and 15 datasets from FinBEN (Xie et al., 2024a): (1) sentiment analysis (FiQA-SA, FOMC, FPB, Headlines); (2) named entity recognition (Finer-Ord, NER); (3) number understanding (ConvFinQA, FinQA); (4) text summarization (ECTSUM, EDTSUM); (5) stock movement prediction (ACL18, BigData22, CIKM18); and (6) credit scoring (Australia, German).
@@
Overall, Table 7 shows that FinLLaMA-Instruct shows that FinLLaMA-Instruct outperforms other financial LLMs on 4 out of 6 financial tasks, including surpassing GPT-4 on 3 tasks, underscoring its effectiveness and applicability in the financial domain. FinLLaMA-Instruct also exceeds the performance of all other specialized financial LLMs, including Palmyra-Fin-70B-32, which is significantly larger, in 4 out of 6 financial tasks. In the numerical understanding task, FinLLaMA-Instruct achieves the best performance with an average accuracy score of 0.69, even outperforming GPT-4. This highlights the effectiveness of our instruction tuning using large-scale math reasoning data in enhancing the model’s numeric understanding ability. Furthermore, FinLLaMA-Instruct consistently outperforms Mistral-7B-Instruct on all tasks and surpasses ChatGPT in five out of six tasks, with text summarization being the exception. This underscores its superior performance compared to general LLMs. FinLLaMA-instruct achieves better performance compared with GPT-4 on three key financial analysis tasks, demonstrating the robustness of the FinLLaMA backbone model and the effectiveness of our fine-tuning approach and datasets.
@@ 3.3 Performance of FinLLaVA
Multimodal Tasks We evaluate our model on four multimodal understanding tasks (Table 11). The MMMU-Overall dataset (10,500 instances) assesses general multimodal capabilities, while MMMU-Business (1,428 instances) focuses on financial domains such as accounting and marketing. In addition to existing datasets, we introduce two new financial multimodal benchmarks. ChartBench tests chart interpretation skills with 350 finance-related instances selected from ChartInstructionData. We used GPT-4o to assign finance-relevance scores and selected samples scoring 9 or above (out of 10), categorized into seven groups (Appendix F.3), with 50 randomly sampled instances per category. TableBench evaluates multimodal reasoning over tabular images, providing a realistic setting for complex financial data interpretation. It contains 450 questions across comparison and data retrieval tasks, reflecting key decision-making processes in finance.
Performance As shown in Table 6, our multimodal model, FinLLaVA, achieves the best performance across all tasks among open-source 7B and 13B models, even surpassing larger models like LLaVA-1.5 and LLaVA-1.6 (Vicuna-13B). On TableBench, FinLLaVA not only outperforms open-source baselines but also surpasses commercial state-of-the-art models such as GPT-4 and Gemini-1.5-pro, demonstrating the effectiveness of our multimodal extension for complex financial data. These results highlight FinLLaVA’s robustness and generalization, especially given that our evaluation uses out-of-domain data, unlike FinTral. By incorporating extensive OCR data during alignment, FinLLaVA improves its ability to interpret financial tables. With fully image-based inputs, users can directly provide table or chart images, making it practical for real-world applications like financial reporting and auditing. This simplicity enables financial professionals to streamline workflows and enhance productivity without technical expertise. Overall, FinLLaVA achieves state-of-the-art performance across multimodal financial tasks, surpassing both open-source and commercial models, with strong generalization and practical utility for real-world applications like financial reporting and auditing.
@@ 4 Conclusion
Appendix A Related Work
A.1 Financial Large Language Models
A.2 Domain Specialization of Large Language Models
Appendix B Curation of Continual Pre-training Corpus
B.1 Dataset Details
@@
B.1.1 Data Processing & Cleaning
@@
B.2 Tabular and Time-series Data Format
@@
Appendix C Financial Instruction Dataset
C.1 Data Sources
@@ C.2 Data Processing
Appendix D Multimodal Instruction Data
D.1 Table
@@
To ensure OCR quality, we limit tables to a maximum size of $10\times 10$ to avoid resolution-related cell blurring. In the SFT stage, we design seven financial-specific tasks (Appendix F.3), randomly sampled in the prompts. Figure 2 shows an example table and how we align and generate SFT data.
D.2 Chart
@@
During SFT, we develop seven chart-specific tasks (Appendix F.3) to enhance the model’s ability to interpret financial charts and extract insights. Figure 4 illustrates an example chart used in SFT data generation.
@@
Appendix E Prompts for Generating Finance-Relevance Score
@@
Appendix F Experiments
F.1 Datasets used for different experiments
@@
F.2 Details of Agent evaluation results
Results of different models in agent evaluation setting is lised in F.2.
@@ F.2 Details of Agent evaluation results
F.3 Multimodal Task Categories
@@
F.4 Details of Continual Pretrained Model Evaluation
F.4.1 Descriptions of Zero-shot Evaluation Tasks
@@
F.4.2 Descriptions of Few-shot Evaluation Tasks
@@
F.5 Prompts for Multimodal Benchmark Evaluation
@@
F.6 Prompts for Generating LLM Answers
@@ F.7 Example of LLM Answers
F.8 Prompts for Evaluating LLM Answers
@@
Appendix G Details for Trading in the Decision Making Task
@@ 4. Outputs:
Appendix H Trading Comparison in the Decision Making Task

### v3

@@ 4.2 Performance of FinLLaMA-Instruct
Table 10 provides detailed information on the datasets and tasks used for the evaluation of instruction fine-tuned models. We align our evaluations with Fintral’s (Bhatia et al., 2024) settings for consistency and comparability, reporting the average performance across datasets in each task. The evaluation includes 6 tasks and 15 datasets used in the FinBEN paper (Xie et al., 2024a): (1) sentiment analysis (SA) task, using FiQA-SA, FOMC, FPB, and Headlines datasets; (2) named entity recognition (NER) task, using Finer-Ord and NER datasets; (3) number understanding (NU) task, using ConvFinQA and FinQA datasets; (4) text summarization (TS) task, using ECTSUM and EDTSUM datasets; (5) stock movement prediction (SMP) task, using ACL18, BigData22, and CIKM18 datasets; (6) credit scoring (CS) task, using Australia and German datasets.
Overall, Table 7 shows that FinLLaMA-Instruct outperforms other financial LLMs on 4 out of 6 financial tasks, including surpassing GPT-4 on 3 tasks, underscoring its effectiveness and applicability in the financial domain. FinLLaMA-Instruct also exceeds the performance of all other specialized financial LLMs, including Palmyra-Fin-70B-32, which is significantly larger, in 4 out of 6 financial tasks. In the numerical understanding task, FinLLaMA-Instruct achieves the best performance with an average accuracy score of 0.69, even outperforming GPT-4. This highlights the effectiveness of our instruction tuning using large-scale math reasoning data in enhancing the model’s numeric understanding ability. Furthermore, FinLLaMA-Instruct consistently outperforms Mistral-7B-Instruct on all tasks and surpasses ChatGPT in five out of six tasks, with text summarization being the exception. This underscores its superior performance compared to general LLMs. FinLLaMA-Instruct achieves better performance compared with GPT-4 on three key financial analysis tasks, demonstrating the robustness of the FinLLaMA backbone model and the effectiveness of our fine-tuning approach and datasets.
@@ 4.3 Performance of FinLLaVA
Multimodal Tasks We evaluate our model on four multimodal understanding tasks as shown in Table 11. The MMMU-Overall dataset, with 10,500 instances, assesses general multimodal capabilities, while the MMMU-Business dataset, with 1,428 instances, evaluates performance in financial domains such as accounting and marketing. In addition to existing datasets, in this paper, we build two new financial multi-modal evaluation tasks. First, ChartBench tests chart interpretation skills with 350 financially relevant instances selected using the ChartInstructionData. We employed GPT-4o to assign a finance-relevance score, selecting instances with scores of 9 or above (out of 10). These were categorized into seven groups, as detailed in Appendix E.3, with 50 randomly chosen instances per category forming the ChartBench benchmark. The TableBench assesses multimodal capabilities using tabular images, offering a realistic testbed for handling complex financial data. Our dataset includes 450 questions split between comparison and data retrieval tasks, essential for extracting data points and comparing metrics, reflecting key decision-making processes in finance.
Performance As shown in Table 6, our multimodal model, FinLLaVA, our multimodal model, FinLLaVA, achieves the best performance across all tasks among open-source models with 7B and 13B model sizes. FinLLaVA even outperforms larger models like LLaVA-1.5 and LLaVA-1.6, both of which use Vicuna-13B. On TableBench, FinLLaVA achieves the best performance and outperforms SOTA commercialized LLMs GPT-4 and Gemini-1.5-pro, which proves the effectiveness of our multimodal extension. These results highlight the robustness and promising performance of FinLLaVA On the TableBench dataset, FinLLaVA not only achieves the best performance but also surpasses state-of-the-art commercialized LLMs like GPT-4 and Gemini-1.5-pro. This success demonstrates the effectiveness of our multimodal extension in enhancing the model’s capability to process and analyze complex financial data. These results highlight its potential for widespread application in the financial domain, offering an efficient solution for interpreting and managing multimodal financial data. Our evaluation differs from FinTral’s in using out-of-domain data, showcasing our model’s robustness and superior performance. We enhance its ability to interpret financial tables by integrating extensive OCR data during alignment. With entirely image-based input, users only need to provide table or chart images, making it convenient for real-world applications like financial reporting and auditing. This simplicity allows financial professionals to efficiently use our model, streamlining their workflow and improving productivity without requiring extensive technical knowledge.
@@ 5 Conclusion
Limitation
Our Open-FinLLMs, including the foundational FinLLaMA, the instruction-tuned FinLLaMA-Instruct, and the multimodal FinLLaVA, demonstrate promising capabilities but have several limitations. First, the models are currently limited to a size of 8B parameters, and future work should explore both smaller models for efficiency and larger models for enhanced performance. Second, our experiments focused exclusively on English, highlighting the need to expand to multilingual settings to better serve global financial markets. Third, the tasks we addressed, such as trading-related scenarios, represent only a subset of potential applications, and future research should investigate broader industrial use cases like financial auditing, risk management, and regulatory compliance. Finally, the multimodal capabilities of FinLLaVA are restricted to handling charts and tabular data; incorporating other data types.
Ethical Statement
Appendix A Curation of Continual Pre-training Corpus
A.1 Dataset Details
@@
A.1.1 Data Processing & Cleaning
@@
A.2 Tabular and Time-series Data Format
@@
Appendix B Financial Instruction Dataset
B.1 Data Sources
@@ B.2 Data Processing
Appendix C Multimodal Instruction Data
C.1 Table
@@
To ensure OCR quality, we limit tables to a maximum size of $10\times 10$ to avoid resolution-related cell blurring. In the SFT stage, we design seven financial-specific tasks (Appendix E.3), randomly sampled in the prompts. Figure 2 shows an example table and how we align and generate SFT data.
C.2 Chart
@@
During SFT, we develop seven chart-specific tasks (Appendix E.3) to enhance the model’s ability to interpret financial charts and extract insights. Figure 4 illustrates an example chart used in SFT data generation.
@@
Appendix D Prompts for Generating Finance-Relevance Score
@@
Appendix E Experiments
E.1 Datasets used for different experiments
@@
E.2 Details of Agent evaluation results
Results of different models in agent evaluation setting is lised in E.2.
@@ E.2 Details of Agent evaluation results
E.3 Multimodal Task Categories
@@
E.4 Details of Continual Pretrained Model Evaluation
E.4.1 Descriptions of Zero-shot Evaluation Tasks
@@
E.4.2 Descriptions of Few-shot Evaluation Tasks
@@
E.5 Prompts for Multimodal Benchmark Evaluation
@@
E.6 Prompts for Generating LLM Answers
@@ E.7 Example of LLM Answers
E.8 Prompts for Evaluating LLM Answers
@@
Appendix F Details for Trading in the Decision Making Task
@@ 4. Outputs:
Appendix G Trading Comparison in the Decision Making Task